There are 640 acres in a square mile and 5280 feet in a mile. A rectangular one-acre lot has perimeter of 836 feet. Find the dimensions of that lot.
Answer by Alan3354(69443)

1 acre = 43560 sq ft
----------------
L*W = 43560
2L + 2W = 836 --> W = 418 - L
Sub for W in eqn 1
L*(418 - L) = 43560
Solve the quadratic